OpenStudy (anonymous):

(4r+5)2>or equal to 6r

OpenStudy (anonymous):

8r + 10 >or equal to 6r -8r -8r 10 >or equal to -2r /2 /2 5 >or equal to -r *-1 *-1 -5<or equal to r

OpenStudy (radar):

8r+10>or=6r 2r>/=-10\[r \ge-5\]

OpenStudy (anonymous):

r<or=-5 thank you

OpenStudy (radar):

I think r has to be equal to or greater than -5.

OpenStudy (anonymous):

when there's a negative number the sign must be reverse

OpenStudy (radar):

We both came up to that conclusion, how did you turn it around so that r is less than or equal -5?

OpenStudy (anonymous):

the inequlatity sign must be reverse

OpenStudy (anonymous):

the -r is less than the 5, so you multiply by -1 to make the positive r greater than -5

OpenStudy (anonymous):

thank you damonky,

OpenStudy (radar):

Only if the equation was multiplied or divided by a negative number. In my solution I divided thru by a POSITIVE 2. To double check, substitute a test number. You say r must be less than -5, so let r=-6 and substitute in orig equation. (4(-6)+5)2 > or equal 6(-6) (-24+5)2 > or = -36 (-19)(2)>or equal -36 -38 > or = -36 Don't think so

OpenStudy (anonymous):

Thanks for the check Radar.....

OpenStudy (anonymous):

From Mathematica:\[\text{Reduce}[(4r+5)2\text{$>$=} 6r] \to r\geq -5 \]